OpenStudy (anonymous):

By 3 pm, Erin had plowed 7 driveways, and by 530pm she had completed 13. a. Find Erin's plowing rate, in number of driveways per hour. b. Find Erin's plowing rate, in minutes per driveway.

OpenStudy (anonymous):

I know the first one will be 13-7/530-3 = 6/2.5 = 2.4 driveways per hour. But I am confused on how to do b.

OpenStudy (anonymous):

2.4 driveways per hour = 2.4/60minutes 60/2.4 = 25 minutes per driveway

OpenStudy (anonymous):

Thanks.